A ball is thrown upward with a speed of 3m per second from the tower of 50m high.find the speed with which it strikes the ground.

Grade:11

1 Answers

Manish Tiwari
6 years ago
Answer:speed=31.5 m/sExplanation:We`re asked to find the speed with which a ball hits the ground, given that it is launched upward from a 50-m tall building with an initial velocity of 3 m/s.To do this, we can use the equation(vy)2−(v0y)2+2ay(y−y0)wherevy is the final y-velocity (what we`re trying to find)v0y is the initial y-velocity (given as positive 3 m/s)ay is the y-acceleration, in this case −g (−9.81 m/s2)y is the final height (ground level, y=0)y0 is the initial height (y=50 m)Plugging in known values, we have(vy)2=(3lm/s)2−2(9.81lm/s2)(0−50lm)=9lm2/s2+981 m2/s2vy=√9lm2/s2+981lm2/s2=±31.5 m/sSince it strikes the ground traveling downward, its velocity then is −31.5 m/s.However, we were asked for the speed of the ball when it strikes the ground. This is merely the magnitude of the velocity:speed=|−31.5lm/s|=31.5 m/sHope it helps!